Roma Sherry drove
360360
miles from her home town to her​ parents' town. During her return​ trip, she was able to increase her speed by
44
miles per hour. If her return trip took 1 hour less​ time, find her original speed and her speed returning home.

Found 2 solutions by jorel555, ikleyn:
Answer by jorel555(1290)

Assuming those numbers are actually doubles, and Roma really drove 360 miles, and increased her speed by 4 mph; then let s be her speed getting there. Then, her speed going back is s+4. So:
360/s=360/(s+4) + 1
360(s+4)=360s+sē+4s
sē+4s-1440=0
(s+40)(s-36)=0
s=36 or -40
Throwing out the negative result, we get her speed as 36 mph on the way there; and 40 mph on the way back. ☺☺☺☺

Answer by ikleyn(52847)

.
When will you learn on how to input your numbers correctly?